### v3

# arXiv:2601.12458v3[math.CA] 28 Feb 2026

arXiv:2601.12458v3[math.CA] 28 Feb 2026
Symmetric Preparation of Systems
Nils Dencker
Lund University
Department of Mathematics, University of Lund, Box 118, S-221 00 Lund, Sweden nils.dencker@gmail.com
1. Introduction and statement of results
If $f$ is $C^{\infty}$ satisfying $\partial_{t}^{m}f(0,0)\neq 0$ and $\partial_{t}^{j}f(0,0)=0$ for $j<m$, then by the Malgrange preparation theorem there exist $C^{\infty}$ functions $c(t,x)$ and $r(t,x)=\sum_{j=0}^{m-1}t^{j}r_{j}(x)$ such that $c(0,0)\neq 0$, $r_{j}(0)=0,\ \forall\,j$ and
(1.1) $f(t,x)=c(t,x)\left(t^{m}+r(t,x)\right)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
locally near $(0,0)$. In the analytic case, the Weierstrass preparation theorem, one has uniqueness but not in the $C^{\infty}$ case, except for $m=1$ and real valued $r(t,x)$ when one can use the implicit function theorem. There are also corresponding division results.
The Malgrange preparation theorem has been generalized to $N\times N$ systems of functions in [3]. An example is for smooth $N\times N$ system $F(t,x)$ satisfying
(1.2) $|\partial_{t}^{m}F(0,0)|>0\quad\text{and}\quad\partial_{t}^{j}F(0,0)=0,\quad j<m$
Then there exist smooth $N\times N$ systems $C(t,x)$ and $M_{j}(x)$, $j<m$, such that $|C(0,0)|\neq 0$, $M_{j}(0)=0$, $\forall\,j$, and
(1.3) $F(t,x)=C(t,x)\left(t^{m}\operatorname{Id}_{N}+\sum_{j=0}^{m-1}t^{j}M_{j}(x)\right)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
locally near $(0,0)$, this is the special case of [3, Th. 5.3]. In the analytic case, one has uniqueness but not in the smooth case.
We also proved the following division theorem in [3]. If $F(t,x)$ is a smooth $N\times N$ system satisfying (1.2), then for any smooth $N\times N$ system $G(t,x)$ there exist smooth $N\times N$ systems $C(t,x)$ and $R_{j}(x)$, $j<m$, such that
(1.4) $G(t,x)=C(t,x)F(t,x)+\sum_{j=0}^{m-1}t^{j}R_{j}(x)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
locally near $(0,0)$, this is a special case of [3, Th. 5.9]. As before, one has uniqueness in the analytic case but not in the smooth case. By taking the adjoints, we also obtain adjoint preparation results and division results, see [3].
Now, the question about symmetric preparation results was raised by Gregory Berkolaiko [1]. The simplest case is for smooth (or analytic) symmetric $N\times N$ matrices $F(t,x)$ satisfying $\partial_{t}F(0,0)>0$ and $F(0,0)=0$. Then the question was: does it exist smooth (or analytic) $N\times N$ matrices $U(t,x)$ and $M(x)$, such that $|U(0,0)|\neq 0$, $M(x)$ is symmetric with $M(0)=0$, satisfying
(1.5) $F(t,x)=U(t,x)\left(t\operatorname{Id}_{N}+M(x)\right)U^{*}(t,x)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
near $(0,0)$, and is this preparation unique? Observe that the conditions on $F(t,x)$ are necessary for (1.5) to hold since $U(0,0)U^{*}(0,0)=\partial_{t}F(0,0)>0$ and $F(0,0)=U(0,0)M(0)U^{*}(0,0)$. Here and in the following we say that the matrix $A$ is symmetric if $A^{*}=A$ and antisymmetric if $A^{*}=-A$, where $A^{*}=\overline{A}^{t}$, and when we say that $A>0$ then we assume that $A$ is symmetric. This result was proved in [4], and in this paper we shall generalize to the higher order case. The following is the main result of the paper.
Theorem 1.1.
Assume that $F(t,x)$ is a smooth symmetric $N\times N$ matrix valued function of $(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$ satisfying
(1.6) $\partial_{t}^{m}F(0,0)>0\quad\text{and}\quad\partial_{t}^{j}F(0,0)=0,\quad j<m$
then there exists a neighborhood $\omega$ of $(0,0)$ and smooth $N\times N$ matrix valued functions $U(t,x)$ and $M_{j}(x)$, $j<m$, such that $|U(0,0)|\neq 0$, $M_{j}(x)$ is symmetric such that $M_{j}(0)=0,\ \forall\,j,$ and
(1.7) $F(t,x)=U(t,x)\left(t^{m}\operatorname{Id}_{N}+\sum_{j=0}^{m-1}t^{j}M_{j}(x)\right)U^{*}(t,x)\quad\text{in $\omega$}$
If $F(t,x)$ is analytic then we obtain (1.7) with analytic $U(t,x)$ and ${\mathbb{M}}(t,x)$. One can choose $U>0$ and then the preparation is unique.
Remark 1.2.
Observe that the preparation (1.7) is not unique in general. For example, if $U(t,x)$ and ${\mathbb{M}}(t,x)$ are solutions to (1.7), then for any orthogonal $A(x)=(A^{-1})^{*}(x)$ we find that $\widetilde{U}(t,x)=U(t,x)A(x)$ and $\widetilde{\mathbb{M}}(t,x)=A^{*}(x){\mathbb{M}}(t,x)A(x)$ solves (1.7). But if $U(t,x)>0$ and $UA(t,x)>0$ then $A(t,x)\equiv\operatorname{Id}_{N}$.
In fact, if $U>0$ then $U$ has a ON base of eigenvectors, and if $UA$ is symmetric then $A^{*}U=UA$ which gives $U=AUA$. Then $A$ preserves the eigenspace with the largest eigenvalue $\|U\|$ and $A^{2}=\operatorname{Id}$ on this space. By iteration over decreasing eigenvalues of $U$ we find $A^{2}=\operatorname{Id}_{N}$. Then $A=A^{-1}=A^{*}$ has eigenvalues $\pm 1$, and if $UA>0$ there cannot be any negative eigenvalues, thus $A=\operatorname{Id}_{N}$.
Now (1.7) gives that $\partial_{t}^{j}F(0,0)=0$ for $j<m$ and $m!U(0,0)U^{*}(0,0)=\partial_{t}^{m}F(0,0)>0$, so the conditions are necessary. If $\partial_{t}^{j}F(0,0)\neq 0$ for $j<m$ then by subtracting suitable terms we obtain the following result.
Corollary 1.3.
Assume that $F(t,x)$ is a smooth symmetric $N\times N$ matrix valued function of $(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$ satisfying
(1.8) $\partial_{t}^{m}F(0,0)>0$
Then for any smooth symmetric $N\times N$ system $\mathbb{R}(t,x)$ such that $\partial_{t}^{j}\mathbb{R}(0,0)=\partial_{t}^{j}F(0,0)$, $j<m$, and $\partial_{t}^{m}\mathbb{R}(0,0)<\partial_{t}^{j}F(0,0)$, there exists a neighborhood $\omega$ of $(0,0)$ and smooth $N\times N$ matrix valued functions $U(t,x)$ and ${\mathbb{M}}(t,x)$ such that $|U(0,0)|\neq 0$, ${\mathbb{M}}(t,x)=\sum_{j=0}^{m-1}t^{j}M_{j}(x)$ is symmetric such that $M_{j}(0)=0,\ \forall\,j$, and
(1.9) $F(t,x)=U(t,x)\left(t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t,x)\right)U^{*}(t,x)+\mathbb{R}(t,x)\qquad\text{in $\omega$}$
If $F(t,x)$ is analytic then we obtain (1.9) with analytic $U(t,x)$, ${\mathbb{M}}(t,x)$ and $\mathbb{R}(t,x)$. One can choose $U>0$ and then the preparation is unique.
Note that the uniqueness does not hold in general, since Remark 1.2 gives counterexamples for nonsymmetric $U$. Since (1.7) is not linear in $U(t,x)$ the proof approaches of Weierstrass or Malgrange preparation results do not work in the symmetric case, instead we shall use inverse function theorems. Theorem 1.1 will be proved in Section 3 for the analytic case, and in Section 4 for the smooth case. But first we shall consider the formal power series case.
2. Formal power series solutions
We shall assume that $F(t,x)$ is given by a formal power series expansion:
(2.1) $F(t,x)=\sum_{{j,k}}F_{j,\alpha}t^{j}x^{k}$
where $F_{j,0}=\partial_{t}^{j}F(0,0)/j!=0$ for $j<m$ and $F_{m,0}=\partial_{t}^{m}F(0,0)/m!>0$. We shall take the formal power series expansions $U(t,x)=\sum_{{j,k}}U_{\begin{subarray}{c}j,k\end{subarray}}t^{j}x^{k}$ where $|U_{0,0}|\neq 0$ and
(2.2) ${\mathbb{M}}(t,x)=\sum_{j<m}t^{j}{\mathbb{M}}_{j}(x)=\sum_{\begin{subarray}{c}j<m\\ k>0\end{subarray}}M_{j,n}t^{j}x^{k}$
is symmetric such that ${\mathbb{M}}_{j}(0)=M_{j,0}=0$, $j<m$.
Then (1.7) gives
(2.3) $\sum_{j,k}F_{j,k}t^{j}x^{k}=\sum_{j,k}U_{j,k}t^{j}x^{k}\Big(t^{m}\operatorname{Id}_{N}+\sum_{\begin{subarray}{c}j,k\end{subarray}}M_{j,k}t^{j}x^{k}\Big)\sum_{{j,k}}U_{j,k}^{*}t^{j}x^{k}$
and by matching coefficients we obtain that
(2.4) $F_{j,k}=\sum_{i,\ell}U_{i,\ell}U_{j-m-i,k-\ell}^{*}+\sum_{\begin{subarray}{c}i,\ell,p,q\end{subarray}}U_{i,\ell}M_{p,q}U_{j-p-i,k-q-\ell}^{*}$
where we sum over the nonnegative index. The first terms of this expansion are $F_{0,0}=U_{0,0}M_{0,0}U_{0,0}^{*}=0$, $F_{0,1}=U_{0,0}M_{0,1}U_{0,0}^{*}$, $F_{j,0}=0$, $j<m$, and $F_{m,0}=U_{0,0}U_{0,0}^{*}$. Here we can choose $U_{0,0}=(F_{m,0})^{1/2}$ which is the unique positive definite symmetric solution. In fact, since $U_{0,0}$ is invertible we can write any other positive symmetric solution as $U_{0,0}A$ and then $(U_{0,0}A)^{*}=A^{*}U_{0,0}$. Since $(U_{0,0}A)^{2}=U_{0,0}^{2}$, $A$ has to be orthogonal. But then Remark 1.2 gives that $A=\operatorname{Id}_{N}$. Thus, we get a unique $U_{0,0}>0$.
When $j<m$ we obtain from (2.4) that
(2.5) $U_{0,0}M_{j,k}U_{0,0}^{*}=F_{j,k}-\sum_{\begin{subarray}{c}i+\ell<j+k\end{subarray}}U_{i,\ell}M_{p,q}U_{j-p-i,k-q-\ell}^{*}$
since $|U_{0,0}|\neq 0$ we obtain symmetric $M_{j,k}$ that is given by $F_{j,k}$, $U_{i,\ell}$ with $i\leq j$ and $\ell\leq k$, and by $M_{p,q}$ with $p\leq j$, $q\leq k$ and $p+q<j+k$, Thus $M_{j,k}$ is determined by $F_{j,k}$, $U_{i,\ell}$ with $\ell<k$ or $\ell=k$ and $i<j$ and by $M_{p,q}$ with $q<k$ or $q=k$ and $p<j$.
With $j$ replaced by $j+m$, we also obtain from (2.4) that
(2.6) $2\operatorname{Re}U_{j,k}U_{0,0}^{*}=F_{j+m,k}-\sum_{\begin{subarray}{c}0<i+\ell<j+k\end{subarray}}U_{i,\ell}U_{j-i,k-\ell}^{*}-\sum_{q>0}U_{i,\ell}M_{p,q}U_{j+m-p-i,k-q-\ell}^{*}$
where the right hand side is given by $F_{j+m,k}$, $U_{i,\ell}$ with $i\leq j+m$ and $\ell<k$ or with $i+\ell<j+k$, thus by $\ell<k$ or $\ell=k$ and $i<j$, and $M_{p,q}$ with $p<m$ and $q\leq k$. Since $U_{0,0}$ is invertible, we can solve (2.6). In the case $U_{0,0}>0$, Lemma A.1 gives a unique symmetric $U_{j,k}$.
We can now determine $U_{j,k}$ and $M_{j,k}$ from (2.5) and (2.6) by induction over $j\geq 0$ for fixed $k\geq 0$ starting with $U_{0,k}$ and $M_{0,k}$ which we first determine by induction over $k$. Since $M_{0,0}=0$ we start with $U_{0,0}$ and $M_{0,1}=U_{0,0}^{-1}F_{0,1}(U_{0,0}^{*})^{-1}$, and assume that $F_{j,k}$ is known. We assume that $U_{0,\ell}$ and $M_{0,\ell+1}$ is known for $\ell<k$. Then $U_{0,k}$ is determined by these and then $M_{0,k+1}$ is determined by $U_{0,\ell}$ and $M_{0,\ell}$ for $\ell\leq k$. By induction we obtain $U_{0,k}$ and $M_{0,k+1}$ for any $k\geq 0$.
Then we use induction over $j\geq 0$ for $U_{j,k}$ and $M_{j,k+1}$ starting with $k=0$. We assume that we have determined $U_{i,\ell}$ and $M_{i,\ell+1}$ when $\ell<k$ and when $\ell=k$ for $i<j$. Then $U_{j,k}$ is determined by these and we then obtain $M_{j,k+1}$ from $U_{i,\ell}$ when $\ell\leq k$ and from $M_{i,\ell+1}$ when $\ell<k$ and when $\ell=k$ for $i<j$. By induction we obtain $U_{j,k}$ and $M_{j,k+1}$ for any $j$ and $k$. Observe that any finite expansion of $U(t,x)$ is positive definite if $|t|+|x|\ll 1$.
Theorem 2.1.
If $F(t,x)$ is analytic (or formal power series) and symmetric such that $\partial_{t}^{m}F(0,0)>0$ and $\partial_{t}^{j}F(0,0)=0$, $\forall\,j<m$, then there exists formal power series $U(t,x)$ and ${\mathbb{M}}(t,x)$ given by (2.2) solving (1.7) modulo $O(|t|^{N}+|x|^{N})$, $\forall\,N$. We get a unique symmetric power series solution in the case when $U_{0,0}>0$.
But is it not clear that in the analytic case, the formal power expansion converges to an analytic function. Instead we will use the inverse function theorem, see Section 3.
3. Proof of Theorem 1.1 in the analytic case
It is clear that $N\times N$ analytic systems form a Banach space with the usual $L^{\infty}$ norm, and they can be localized to neighborhoods. We shall use the Taylor expansion notation
$T_{k}(F)(t,x)=\sum_{j=0}^{k-1}t^{j}\partial_{j}^{j}F(0,x)/j!$
and also the $k$ times integrations $I_{k}(F)(t,x)$ in $t$ such that $I_{1}(F)(t,x)=\int_{0}^{t}F(s,x)\,ds$ and $I_{k+1}(F)(t,x)=I_{1}(I_{k}(F))(t,x)$. Thus we find $\partial_{t}^{j}I_{k}(F)(0,x)\equiv 0$ when $j<k$ and $\partial_{t}^{k}I_{k}(F)(t,x)\equiv F(t,x)$. Observe that $t^{-k}I_{k}(F)(t,x)$ is analytic or smooth if $F(t,x)$ is.
We shall prove that if $U(t,x)$ and ${\mathbb{M}}(t,x)$ are analytic such that $|U(0,0)|\neq 0$, ${\mathbb{M}}(t,x)=\sum_{j=0}^{m-1}t^{j}M_{j}(x)/j!$ is symmetric such that $M_{j}(0)=0,\ \forall\,j$, then the mapping
(3.1) $(U(t,x),{\mathbb{M}}(t,x))\overset{\mathcal{F}}{\longrightarrow}\\ \Big(\operatorname{Re}\partial_{t}^{m}\big(U(t,x)(t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t,x))U^{*}(t,x)\big),T_{m}(U{\mathbb{M}}U^{*})(t,x))\Big)\\ =(F_{1}(t,x),F_{0}(t,x))$
is analytic and surjective in some neighborhood of $(0,0)$ for analytic and symmetric $F_{j}$, j=0, 1j=0,\,1, satisfying $|F_{1}(0,0)|\neq 0$, $\partial_{t}^{m}F_{0}\equiv 0$ and $F_{0}(t,0)\equiv 0$. If $\mathcal{F}(U,{\mathbb{M}})=(F_{1},F_{0})$ by (3.1) in a neighborhood of $(0,0)$ then we obtain (1.7) with
(3.2) $F(t,x)=F_{0}(t,x)+I_{m}(F_{1})(t,x)\qquad\text{near $(0,0)$}$
since $F_{0}(t,x)=T_{m}F(t,x)$ and $F_{1}(t,x)=\partial_{t}^{m}F(t,x)$.
When ${\mathbb{M}}\equiv 0$ we find that $F_{0}\equiv 0$ and $\partial_{t}^{m}(t^{m}UU^{*})=m!UU^{*}=F_{1}$ when $t=0$ so $U(t,x)$ must be invertible close to $(0,0)$. For any such $U(t,x)$, we find that the differential of $\mathcal{F}$ at $(U,0)$ is given by
(3.3) $({\bf u}(t,x),{{\bf m}}(x))\overset{d\mathcal{F}}{\longrightarrow}\\ \Big(\partial_{t}^{m}\big(2t^{m}\operatorname{Re}U(t,x){\bf u}^{*}(t,x)+U(t,x){{\bf m}}(t,x)U^{*}(t,x)\big),T_{m}(U{\bf m}U^{*})(t,x)\Big)\\ =(A_{1}(t,x),A_{0}(t,x))$
where ${\bf u}(t,x)$ and ${\bf m}(t,x)$ are analytic such that $\bf m$} is symmetric, $\partial_{t}^{m}{\bf m}\equiv 0$ and ${\bf m}(t,0)\equiv 0$, which gives analytic symmetric $A_{j}$, j=0, 1j=0,\,1, such that $\partial_{t}^{m}A_{0}\equiv 0$ and $A_{0}(t,0)\equiv 0$.
We shall show that this differential is surjective for such systems in a neighborhood of $(0,0)$. We find from (3.3) that ${\bf m}=T_{m}(U^{-1}A_{0}(U^{-1})^{*})$ and
$2t^{m}\operatorname{Re}U(t,x){\bf u}^{*}(t,x)=I_{m}(A_{1})(t,x)+\mathbb{R}(t,x)-U{\bf m}U^{*}(t,x)$
for some analytic symmetric $\mathbb{R}(t,x)$ such that $\partial_{t}^{m}\mathbb{R}\equiv 0$. Since the right hand side has to vanish of order $m$ at $t=0$, we find that $\mathbb{R}(t,x)=T_{m}(U{\bf m}U^{*})(t,x)=A_{0}(t,x)$ and then
$2\operatorname{Re}U(t,x){\bf u}^{*}(t,x)=t^{-m}\Big(I_{m}(A_{1})(t,x)+A_{0}(t,x)-U{\bf m}U^{*}(t,x)\Big)=A(t,x)$
which is analytic and symmetric. By substituting $\bf m$} in $A$ we find that $A$ depends linearly on $A_{j}$, j=0, 1j=0,\,1, and that there exists an analytic and antisymmetric $B(t,x)$ so that $2U{\bf u}^{*}=A+B$, which gives $u^{*}=U^{-1}(A+B)/2$ and the solution
(3.4) ${\bf u}(t,x)=\frac{1}{2}(A(t,x)-B(t,x))(U^{-1})^{*}(t,x)$
This gives the surjectivity of $d\mathcal{F}$ in (3.3) and by the inverse function theorem we obtain local surjectivity of the nonlinear mapping $\mathcal{F}$ in (3.1) in a neighborhood $V$ of $(U,0)$. Observe that $\mathcal{F}$ will also be surjective on some $(F_{1},F_{0})$ such that $|F_{0}(t,0)|\ll 1$.
For any symmetric ${\mathbb{M}}(t,x)$ such that $\partial_{t}^{m}{\mathbb{M}}(t,x)\equiv 0$ and ${\mathbb{M}}(t,0)=0$, there exists a neighborhood $\omega$ of $(0,0)$ such that $(U,{\mathbb{M}})\in V$ at $\omega$, and then (3.4) gives a solution to (3.1) in $\omega$. But (3.3) is not injective, since any antisymmetric $B$ gives a different solution. Actually, we have for fixed $U$ and ${\mathbb{M}}\equiv 0$ that the mapping $(A_{0},A_{1},B)\mapsto({\bf u},{\bf m})$ is bijective near $(0,0)$.
It remains to prove uniqueness in the case $U(t,x)$ is positive definite near $(0,0)$, which we obtain when $U(0,0)=(F_{1}(0,0)/m!)^{1/2}>0$. Since $\bf m$} already is determined, we shall prove that there is only one symmetric solution ${\bf u}(t,x)$ to equation (3.4). If $\operatorname{Im}u^{*}=0$ then we obtain from Lemma A.1 a unique analytic antisymmetric $B(t,x)$ that solves
$\operatorname{Im}(U^{-1}B)=-\operatorname{Im}(U^{-1}A)\qquad\text{near~$(0,0)$}$
for the analytic $A(t,x)$ in (3.4). Thus the unique analytic symmetric solution is given by
${\bf u}(t,x)=\frac{1}{2}U^{-1}(t,x)(A(t,x)+B(t,x))$
We find that $d\mathcal{F}$ is bijective which gives that $\mathcal{F}$ is a analytic bijection near $(0,0)$ in the analytic symmetric case. ∎
4. Proof of Theorem 1.1 in the smooth case
The smooth $N\times N$ systems form a local Fréchet spaces with the $C^{\infty}(\overline{\omega})$ norms, which are tame if $\overline{\omega}$ is a compact neighborhood with boundary, see Theorem B.8. For tame Fréchet spaces one can use the Nash–Moser theorem, which is the inverse function theorem for tame Fréchet spaces. The difference to the analytic case is that is it not enought to check the local surjectivity of the linearized map when ${\mathbb{M}}(0)\equiv 0$, instead one has to check that for any smooth $U$ and $\mathbb{M}$} in a neighborhood, see Theorem B.12.
So we shall show that if $U(t,x)\in C^{\infty}$, $|U(0,0)|\neq 0$, ${\mathbb{M}}(t,x)=\sum_{j=0}^{m-1}t^{j}M_{j}(x)\in C^{\infty}$ is symmetric such that $M_{j}(0)=0,\ \forall\,j$, then the following $C^{\infty}$ mapping is surjective near $(0,0)$
(4.1) $C^{\infty}\ni(U(t,x),{\mathbb{M}}(t,x))\overset{\mathcal{F}}{\longrightarrow}\\ \Big(\partial_{t}^{m}\big(U(t,x)(t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t,x))U^{*}(t,x)\big),T_{m}(U{\mathbb{M}}U^{*})(t,x))\Big)\\ =(F_{1}(t,x),F_{0}(t,x))\in C^{\infty}$
Here $F_{j}$ is symmetric, j=0, 1,j=0,\,1, $F_{1}(0,0)>0$, $\partial_{t}^{m}F_{0}\equiv 0$ and $F_{0}(t,0)\equiv 0$. As before, if we obtain a solution $(F_{1},F_{0})$ to (4.1) then we obtain $F(t,x)$ from (3.2) so that $F_{0}(t,x)=T_{m}(F)(t,x)$ and $F_{1}(t,x)=\partial_{t}^{m}F(t,x)$. It follows from Example B.4 that the mapping $\mathcal{F}$ is tame of degree $m$.
Now, for any such $(F_{1}(t,x),F_{0}(t,x))$ we may choose $(U_{0}(t,x),{\mathbb{M}}_{0}(t,x))$ so that $U_{0}(t,x)$ is invertible and $U_{0}(0,0)U^{*}_{0}(0,0)\equiv F_{1}(0,0)$ and ${\mathbb{M}}_{0}(t,x)$ is a symmetric polynomial in $t$ of degree less than $m$ such that ${\mathbb{M}}_{0}(t,0)\equiv 0$. Observe that we may choose symmetric $U_{0}$ such that $U_{0}(0,0)=F_{1}^{1/2}(0,0)>0$, and then $U_{0}>0$ near the origin. Since we shall only need local solutions, we may assume that $U_{0}(t,x)$ and ${\mathbb{M}}_{0}(t,x)\in C^{\infty}(\overline{\omega})$ where $\overline{\omega}$ is compact neighborhood of the origin so that $U(t,x)$ is invertible and ${\mathbb{M}}(t,x)$ is symmetric such that $\partial_{t}^{m}{\mathbb{M}}(t,x)\equiv 0$ and ${\mathbb{M}}(t,0)\equiv 0$.
Then we let $V$ be a bounded neighborhood of $(U_{0}(t,x),{\mathbb{M}}_{0}(t,x))$ in $C^{\infty}(\overline{\omega})$ so that the first component is invertible and the second component is symmetric polynomial in $t$ of degree less than $m$. If $U_{0}>0$ then $V$ is a neighborhood with symmetric matrices. Since ${\mathbb{M}}_{0}(t,0)\equiv 0$ we may for $\varepsilon>0$ choose $V$ and a compact neighborhood $\overline{\omega}$ so that $(U,{\mathbb{M}})\in V$ gives $\|{\mathbb{M}}\|\leq\varepsilon$ in $\overline{\omega}$.
We find for any $(U(t,x),{\mathbb{M}}(t,x))\in V$ that the differential of $\mathcal{F}$ at $(U(t,x),{\mathbb{M}}(t,x))$ is equal to
(4.2) $C^{\infty}\ni({\bf u}(t,x),{\bf m}(t,x))\overset{d\mathcal{F}}{\longrightarrow}\\ \Big(\partial_{t}^{m}\big(2\operatorname{Re}(U(t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t,x)){\bf u}^{*})+U{\bf m}U^{*}\big),T_{m}(U{\bf m}U^{*})+2\operatorname{Re}T_{m}(U{\mathbb{M}}{\bf u}^{*})\Big)\\ =(A_{1}(t,x),A_{0}(t,x))\in C^{\infty}$
where ${\bf u}(t,x)$ and ${\bf m}(t,x)$ are smooth such that $\bf m$} is symmetric and $\partial_{t}^{m}{\bf m}\equiv 0$ which gives smooth symmetric $A_{j}$, j=0, 1j=0,\,1, such that $\partial_{t}^{m}A_{0}\equiv 0$.
Next, we shall consider the special case when the functions are constant in $x$. We shall show that if ${\bf u}(t)$ and ${\bf m}(t)\in C^{\infty}(\overline{\omega})$ for a sufficiently small neighborhoods $\overline{\omega}$ of the origin in $\mathbf{R}$, then the mapping $d\mathcal{F}$ is tame and surjective with a tame right inverse. For that, the condition that $\|{\mathbb{M}}\|\ll 1$ in $\overline{\omega}$ is necessary.
Proposition 4.1.
Let $U(t)$ and ${\mathbb{M}}(t)$ be smooth and $N\times N$ valued functions such that $|U(t)|\neq 0$, $\partial^{m}_{t}{\mathbb{M}}(t)\equiv 0$ and $\|{\mathbb{M}}(t)\|\ll 1$. Then in a sufficiently small compact neighborhood $\overline{\omega}$ of the origin, the mapping $d\mathcal{F}$ given by (4.2) is surjective, having a tame local right inverse of degree 3 in $\overline{\omega}$. If $U(t)$ is positive definite, then $d\mathcal{F}$ is bijective and has a tame local inverse of degree 3.
By Theorem B.12, we obtain that the $C^{\infty}$ tame mapping $\mathcal{F}$ is surjective on a neighborhood $W$ of a given $(F_{1}(t),F_{0}(t))$ close to the origin, having a smooth tame right inverse $\mathcal{F}^{-1}$ giving $(U(t),{\mathbb{M}}(t))=\mathcal{F}^{-1}(F_{1}(t),F_{0}(t))$. Now, if $F_{j}(t)=F_{j}(t,0)$ then for $|x|\ll 1$ we obtain that $(F_{1}(t,x),F_{0}(t,x))\in W$, so we obtain the symmetric preparation (1.7) with
(4.3) $(U(t,x),{\mathbb{M}}(t,x))=\mathcal{F}^{-1}(F_{1}(t,x),F_{0}(t,x))$
in a neighborhood of $(0,0)$.
If $U(t)$ is positive definite, then by Theorem B.12 we obtain that the mapping $\mathcal{F}$ is a bijection in a neighborhood $W$ of a given $(F_{1}(t),F_{0}(t))$ near the origin with the tame inverse $\mathcal{F}^{-1}$ giving a unique $(U(t),{\mathbb{M}}(t))=\mathcal{F}^{-1}(F_{1}(t),F_{0}(t))$. If $F_{j}(t)=F_{j}(t,0)$ and $|x|\ll 1$ then we obtain from (4.3) the unique symmetric solution $(U(t,x),{\mathbb{M}}(t,x))$ to (1.7) near the origin which finishes the proof of Theorem 1.1 in the smooth case. ∎
It remains to prove Proposition 4.1. We find from (4.2) that
(4.4) ${\bf m}=T_{m}\left(U^{-1}\big(A_{0}-2\operatorname{Re}(U{\mathbb{M}}{\bf u}^{*})\big)(U^{-1})^{*}\right)$
and if $P(t,{\mathbb{M}}(t))=t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t)$ with symmetric ${\mathbb{M}}(t)=\sum_{j<m}t^{j}M_{j}$ we have
(4.5) $2\operatorname{Re}UP(t,{\mathbb{M}}(t)){\bf u}^{*}=I_{m}(A_{1})(t)-U(t){\bf m}(t)U^{*}(t)+\mathbb{R}(t)=A(t)$
for some smooth symmetric $\mathbb{R}(t)$ such that $\partial_{t}^{m}\mathbb{R}(t)\equiv 0$. Observe that $A$ depends linearly on $\bf m$} by (4.4) and thus linearly on $T_{m}({\mathbb{M}}{\bf u}^{*})$, and that we may assume that $\|{\mathbb{M}}\|=\sum_{j<m}\|M_{j}\|\ll 1$. By (4.5) there exists a smooth antisymmetric $B(t)$ so that
(4.6) $2P(t,{\mathbb{M}}(t)){\bf u}^{*}(t)=U^{-1}(t)I_{m}(A_{1})(t)-{\bf m}(t)U^{*}(t)+U^{-1}(t)\mathbb{R}(t)+U^{-1}(t)B(t)$
We shall keep $B$ fixed in the following, and we may take $B(t)\equiv 0$. Since $\mathbb{R}(t)$ is symmetric and $B(t)$ is antisymmetric, we may replace $\mathbb{R}(t)$ with $\mathbb{R}(t)-T_{m}(B)(t)$ and let $T_{m}(B)\equiv 0$. Then we can replace $B(t)$ by $t^{m}B(t)$ which gives
(4.7) $U^{-1}(t)t^{m}B(t)=P(t,{\mathbb{M}}(t))U^{-1}(t)B(t)-{\mathbb{M}}(t)U^{-1}(t)B(t)$
We shall use the following result, which is the adjoint version of Proposition 3.2 in [3] in the special case when $\pi_{m}=\operatorname{Id}_{N}$.
Proposition 4.2.
Let $G(t)\in\mathcal{S}(\mathbf{R})$ be $N\times N$ valued and let $P(t,{\mathbb{M}}(t))=t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t)$ with $N\times N$ system ${\mathbb{M}}(t)=\sum_{j<m}t^{j}M_{j}$, ${\mathbb{M}}=\left\{\,M_{j}\,\right\}_{j<m}$ and $\|{\mathbb{M}}\|=\sum_{j<m}\|M_{j}\|$. Then we can find $N\times N$ systems $Q(t,{\mathbb{M}},G)\in C^{\infty}$ and $\mathbb{R}(t,{\mathbb{M}},G)=\sum_{j<m}t^{j}R_{j}({\mathbb{M}},G)\in C^{\infty}$ depending linearly on $G(t)$, such that
(4.8) $G(t)=P(t,{\mathbb{M}}(t))Q(t,{\mathbb{M}},G)+\mathbb{R}(t,{\mathbb{M}},G){\quad}\text{if}{\quad}\|{\mathbb{M}}\|<1$
Also, we have the following estimates when $\|{\mathbb{M}}\|<1$
(4.9) $\|\partial_{t}^{{\alpha}}\partial_{{\mathbb{M}}}^{{\beta}}Q(t,{\mathbb{M}},G)\|\leq C_{{\alpha}{\beta}}\int(\|G\|+\|G^{(j)}\|)\,dt{\quad}j={\alpha}+3+m(|{\beta}|+1)$
$\|\partial_{{\mathbb{M}}}^{{\beta}}R_{k}(t,{\mathbb{M}},G)\|\leq C_{{\beta}}\int(\|G\|+\|G^{(j)}\|)\,dt{\quad}j=2+m(|{\beta}|+1)\qquad k<m$
We find that (4.8) gives a smooth surjective tame map $C^{\infty}\ni(Q,\mathbb{R})\mapsto G\in\mathcal{S}$ of degree 0 having a smooth tame right inverse of degree $3+m$, see Example B.14. Observe that the estimate (4.8) is not tame in the parameters $\mathbb{M}$}. Thus, for $\|{\mathbb{M}}\|\ll 1$ we shall use Proposition 4.2 on every term on the right hand side of (4.6) using (4.7), after cut off so that they have compact support. We will denote by $\mathbb{R}_{k}(t)$ the terms that are polynomials of degree less than $m$ in $t$.
First, if $P=P(t,{\mathbb{M}}(t))$ then Proposition 4.2 gives smooth $C_{1}(t)$ and $\mathbb{R}_{1}(t)$ so that
(4.10) $U^{-1}(t)I_{m}(A_{1})(t)=PC_{1}(t)+\mathbb{R}_{1}(t)\qquad\text{near $t=0$}$
where $C_{1}$ and $\mathbb{R}_{1}$ are smooth depending linearly on $\bf m$} and thus linearly on $T_{m}({\mathbb{M}}{\bf u}^{*})$ by (4.4), and that
(4.11) ${\bf m}(t)U^{*}(t)=PC_{2}(t)+\mathbb{R}_{2}(t)\qquad\text{near $t=0$}$
where $C_{2}$ and $\mathbb{R}_{2}$ are smooth and depending linearly on $T_{m}({\mathbb{M}}{\bf u}^{*})$. We also find
(4.12) $U^{-1}(t)=PC_{3}(t)+\mathbb{R}_{3}(t)\qquad\text{near $t=0$}$
with smooth $C_{3}$ and $\mathbb{R}_{3}$, thus $U^{-1}\mathbb{R}=PC_{3}\mathbb{R}+\mathbb{R}_{3}\mathbb{R}$. By (4.9), $C_{3}$ is uniformly bounded when $\|{\mathbb{M}}\|<1$. Thus $\mathbb{R}_{3}(0)$ is invertible if $\|{\mathbb{M}}(0)\|\ll 1$ since $\mathbb{R}_{3}(0)=U^{-1}(0)-{\mathbb{M}}(0)C_{3}(0)$. Finally, we obtain smooth $C_{4}(t)$ and $\mathbb{R}_{4}(t)$ so that
[…]
where $\|C_{4}\|_{k}$ and $\|\mathbb{R}_{4}\|_{k}$ are $\mathcal{O}(\|{\mathbb{M}}\|)$ for any local $C^{k}$ norm. We find from Proposition 4.2 that $C_{j}(t)$ and $\mathbb{R}_{j}(t)$, $\forall\,j$, depend smoothly and tamely of degree $3+m$ on $U(t)$ and ${\mathbb{M}}(t)$.
Summing up, we find from (4.7), and (4.10)–(4.13) that
(4.14) $2Pu^{*}(t)=P\left(C_{1}+C_{2}+C_{3}\mathbb{R}+C_{4}B+U^{-1}B\right)+\mathbb{R}_{1}+\mathbb{R}_{2}+\mathbb{R}_{3}\mathbb{R}+\mathbb{R}_{4}B=PC_{0}+\mathbb{R}_{0}$
locally near the origin in $\mathbb{R}$. Here $C_{0}$ and $\mathbb{R}_{0}$ are smooth depending linearly on $\mathbb{R}$, $B$ and $T_{m}({\mathbb{M}}{\bf u}^{*})$. Thus for fixed $U$ and $\mathbb{M}$} we can write
(4.15) $\left\{\begin{aligned} &C_{0}=C_{00}+C_{01}T_{m}({\mathbb{M}}{\bf u}^{*})+C_{3}\mathbb{R}+C_{4}B+U^{-1}B\\ &\mathbb{R}_{0}=\mathbb{R}_{00}+\mathbb{R}_{01}T_{m}({\mathbb{M}}{\bf u}^{*})+\mathbb{R}_{3}\mathbb{R}+\mathbb{R}_{4}B\end{aligned}\right$.
where $\mathbb{R}_{3}(t)$ is invertible near $t=0$ and $\|C_{4}\|_{k}+\|R_{4}\|_{k}=\mathcal{O}(\|{\mathbb{M}}\|)$, $\forall\,k$. Observe that the products are matrix multiplications, except $C_{01}T_{m}({\mathbb{M}}{\bf u}^{*})$ and $R_{01}T_{m}({\mathbb{M}}{\bf u}^{*})$, which represent linear maps.
Now, since $T_{m}({\mathbb{M}}{\bf u}^{*})$ appears in the equations we shall put $G(t)=T_{m}(G)(t)+t^{m}G_{1}(t)$ for $G\in C^{\infty}$ and $C_{jk}(t)=T_{m}(C_{jk})(t)+t^{m}C_{jk}^{1}(t)$, $\forall\,jk$. Then for fixed $U$, $\mathbb{M}$} and $B$ we obtain that $2{\bf u}^{*}\equiv C_{0}$ and $\mathbb{R}_{0}\equiv 0$ if
(4.16) $\left\{\begin{aligned} &2T_{m}({\bf u}^{*})-T_{m}(C_{01}T_{m}({\mathbb{M}}{\bf u}^{*}))-T_{m}(C_{3}\mathbb{R})\equiv T_{m}(C_{00})+T_{m}(C_{4}B)+T_{m}(U^{-1}B)\\ &-C_{01}^{1}T_{m}({\mathbb{M}}{\bf u}^{*})+2{\bf u}_{1}^{*}-(C_{3}\mathbb{R})_{1}\equiv C_{00}^{1}+(C_{4}B)_{1}+(U^{-1}B)_{1}\\ &\mathbb{R}_{01}T_{m}({\mathbb{M}}{\bf u}^{*})+\mathbb{R}_{3}\mathbb{R}\equiv-\mathbb{R}_{00}-\mathbb{R}_{4}B\end{aligned}\right$.
Observe that the first equation is modulo $\mathcal{O}(t^{m})$, and that
(4.17) $\|T_{m}({\mathbb{M}}{\bf u}^{*})\|_{k}\lesssim\|{\mathbb{M}}\|\|{\bf u}^{*}\|_{k+m-1}$
since $\|{\mathbb{M}}\|_{k}\lesssim\|{\mathbb{M}}\|$ and $\|T_{m}{\bf u}^{*}\|_{k}\lesssim\|{\bf u}^{*}\|_{k+m-1}$ for any local $C^{k}$ norm. Then we get from (4.17) that (4.18) is given by
(4.18) $\begin{pmatrix}2\operatorname{Id}_{N}&0&-T_{m}(C_{3}\ \boldsymbol{\cdot})\\ 0&2\operatorname{Id}_{N}&-(C_{3}\ \boldsymbol{\cdot})_{1}\\ 0&0&\mathbb{R}_{3}\end{pmatrix}$
modulo terms having local $C^{k}$ norms that are $\mathcal{O}(\|{\mathbb{M}}\|)$. Here $\operatorname{Im}(C\ \boldsymbol{\cdot})$ is the mapping $E\mapsto\operatorname{Im}(CE)$ and $T_{m}(C\ \boldsymbol{\cdot})$ is the mapping $E\mapsto T_{m}(CE)$. The system (4.18) is invertible close to $t=0$ since $\mathbb{R}_{3}(0)$ is invertible. Thus, we find that the system (4.16) is invertible in a compact neighborhood $\overline{\omega}$ of the origin if $\|{\mathbb{M}}\|\ll 1$, which can be obtained by shrinking $V$ and $\overline{\omega}$. Since this system is linear, smooth and invertible, we find from Example B.4 that it has a tame smooth inverse of degree 0 which gives a solution $(T_{m}({\bf u}^{*}),{\bf u}^{*}_{1},\mathbb{R})$.
Thus, we obtain a smooth solution ${\bf u}^{*}(t)\equiv T_{m}({\bf u}^{*})(t)+t^{m}{\bf u}^{*}_{1}(t)\equiv C_{0}(t)/2$ for $\mathbb{R}_{0}(t)\equiv 0$ in a sufficiently small compact neighborhood $\overline{\omega}$ of the origin. Observe that the neighborhood $\overline{\omega}$ only depends on the bounds on $U$, $\mathbb{M}$} and $B$. Having found ${\bf u}^{*}(t)$ satisfying (4.5) we obtain $({\bf u}(t),{\bf m}(t))$ by (4.4) in $\overline{\omega}$. This gives surjectivity and a tame local right inverse of degree $3+m$ to $d\mathcal{F}$ in $\overline{\omega}$ for any $(U(t),{\mathbb{M}}(t))\in V$. Observe that different antisymmetric $B$ gives different inverses.
It remains to prove uniqueness in the symmetric case for $(U,{\mathbb{M}})\in V$, when $U$ and ${\bf u}(t)$ are symmetric and $U>0$ which then holds in $V$. We shall show that there is a unique symmetric solution to the equation, and find the corresponding $B(t)$ in (4.6). By adding the equation $\operatorname{Im}{\bf u}(t)\equiv\operatorname{Im}(C_{0})(t)\equiv 0$ with $C_{0}(t)$ given by (4.15) we find as before
(4.19) $\left\{\begin{aligned} &2T_{m}({\bf u})-T_{m}(C_{01}T_{m}({\mathbb{M}}{\bf u}^{*}))-T_{m}(C_{3}\mathbb{R})-T_{m}(C_{4}B)+T_{m}(U^{-1}B)\equiv T_{m}(C_{00})\\ &-C_{01}^{1}T_{m}({\mathbb{M}}{\bf u}^{*})+2{\bf u}_{1}-(C_{3}\mathbb{R})_{1}-(C_{4}B)_{1}-(U^{-1}B)_{1}\equiv C_{00}^{1}\\ &\mathbb{R}_{01}T_{m}({\mathbb{M}}{\bf u}^{*})+\mathbb{R}_{3}\mathbb{R}+\mathbb{R}_{4}B\equiv-\mathbb{R}_{00}\\ &\operatorname{Im}(C_{01}T_{m}({\mathbb{M}}{\bf u}^{*}))+\operatorname{Im}(C_{3}\mathbb{R})+\operatorname{Im}(C_{4}B)+\operatorname{Im}(U^{-1}B)\equiv-\operatorname{Im}(C_{00})\end{aligned}\right$.
which is invertible near $t=0$ when $\|{\mathbb{M}}\|\ll 1$. In fact, the matrix of the system is
(4.20) $\begin{pmatrix}2\operatorname{Id}_{N}&0&-T_{m}(C_{3}\ \boldsymbol{\cdot})&-T_{m}(U^{-1}\ \boldsymbol{\cdot})\\ 0&2\operatorname{Id}_{N}&-(C_{3}\ \boldsymbol{\cdot})_{1}&-(U^{-1}\ \boldsymbol{\cdot})_{1}\\ 0&0&\mathbb{R}_{3}&0\\ 0&0&\operatorname{Im}(C_{3}\ \boldsymbol{\cdot})&\operatorname{Im}(U^{-1}\ \boldsymbol{\cdot})\end{pmatrix}$
modulo terms having local $C^{k}$ norms that are $\mathcal{O}(\|{\mathbb{M}}\|)$. Here $\operatorname{Im}(U^{-1}\ \boldsymbol{\cdot})$ is invertible on antisymmetric matrices by Lemma A.1. Since $\mathbb{R}_{3}(0)$ also is invertible, we find that (4.20) is invertible near the origin. As before, we find that (4.19) is invertible in a sufficiently small compact neighborhood $\overline{\omega}$ of the origin if $\|{\mathbb{M}}\|\ll 1$, which can be obtained by shrinking $V$ and $\overline{\omega}$. Thus, we obtain a unique smooth solution $({\bf u}(t),\mathbb{R}(t),B(t))$ to (4.19) with symmetric $\bf u$}, and by (4.4) a unique smooth symmetric solution $({\bf u}(t),{\bf m}(t))$ to (4.2).
Thus, $d\mathcal{F}$ is a tame bijection in a compact neighborhood $\overline{\omega}$ of the origin for any symmetric $(U(t),{\mathbb{M}}(t))\in V$. Observe that this gives a tame inverse $d\mathcal{F}^{-1}$ of degree $3+m$ by Example B.14, and the neighborhood $\overline{\omega}$ only depends on the bounds on $U$ and $\mathbb{M}$}. ∎
Appendix A A symmetry lemma
In this section, we are going to prove the following result about symmetric and antisymmetric invariance of matrices.
Lemma A.1.
If $U(x)>0$ is a smooth $N\times N$ matrix, then $A(x)\mapsto\operatorname{Re}(UA)(x)$ is a bijection of smooth symmetric matrices and $A(x)\mapsto\operatorname{Im}(UA)(x)$ is a bijection of smooth antisymmetric matrices. If $U(x)$ is analytic, then this holds for analytic $A(x)$.
Proof.
First we consider the case of constant matrices. To simplify, we define the symmetric product $S(U,A)=\frac{1}{2}(UA+AU)$, which is symmetric when $A$ is symmetric and antisymmetric when $A$ is antisymmetric, thus $A\mapsto S(U,A)$ preserves the symmetry. We also have that linear ON mappings $O$ preserves the symmetry, since $OAO^{-1}=OAO^{*}$.
Next, we are going to prove that $A\mapsto S(U,A)$ is injective for both symmmetric and antisymmetric $A$. Since any $N\times N$ matrix can be written uniquely as a sum of a symmetric and an antisymmetric matrix, we will then find that $A\to S(U,A)$ is a bijection for any $N\times N$ matrix $A$. Thus, by using an ON base of eigenvectors to $U$, we find that $U=\{\mu_{j}\delta_{jk}\}_{jk}$ with $\mu_{j}>0,\ \forall\,j$, and $A=\{\alpha_{jk}\}_{jk}$. If $A$ is symmetric then $\alpha_{kj}=\alpha_{jk}$ $\forall\,jk$, and if $A$ is antisymmetric then $\alpha_{kj}=-\alpha_{jk}$ $\forall\,jk$. In both cases, we find that
$S(U,A)=\{(\mu_{j}+\mu_{k})\alpha_{jk}/2\}_{jk}=0$
if and only if $\alpha_{jk}=0$ $\forall\,jk$. So $A\mapsto S(U,A)$ is injective, and thus bijective in both cases. Observe that $|\alpha_{jk}|\leq\|U^{-1}\|\|S(U,A)\|$ since $\mu_{j}^{-1}\leq\|U^{-1}\|$, $\forall\,j$.
Next, we consider the equation
$S(U,A)(x)=S(U(x),A(x))=B(x)\in C^{\infty}$
which has a unique solution $A(x)=\{\alpha_{jk}\}_{jk}(x)$ such that $|\alpha_{jk}(x)|\leq\|U^{-1}(x)\|\|B(x)\|$. By compactness we find for any $x_{0}$ and any sequence $x_{j}\to x_{0}$ that there exixts a subsequence having a limit: $A(x_{j})\to A_{0}$. Then by continuity and uniqueness we find that $\lim_{x\to x_{0}}A(x)=A_{0}=A(x_{0})$, which proves continuity of $A(x)$.
To show that $A(x)$ is differentiable, we put $\Delta_{h,y}A(x)=\frac{A(x+hy)-A(x)}{h}$ for any $x,\,y\in\mathbf{R}^{n}$ and $h>0$. Then we find
(A.1) $\Delta_{h,y}S(U,A)(x)=S(\Delta_{h,y}U(x),A(x+hy))+S(U(x),\Delta_{h,y}A(x))=\Delta_{h,y}B(x)$
which gives that
(A.2) $S(U(x),\Delta_{h,y}A(x))=\Delta_{h,y}B(x)-S(\Delta_{h,y}U(x),A(x+hy))$
where the right hand side is continuous in $h$. As before, we that the left hand side has the unique limit $S(U(x_{0}),y\cdot\partial A(x_{0}))$ for any $y$, giving $S(U,\partial A)=\partial B-S(\partial U,A)$.
Let $A_{\alpha}(x)=\partial^{\alpha}A(x)$ and $A_{k}(x)=\{A_{\alpha}(x)\}_{|\alpha|\leq k}$ and similarly for $U(x)$ and $B(x)$. By induction, we can assume that $A(x)\in C^{k}$ and that
(A.3) $B_{\gamma}(x)=\sum_{\alpha+\beta=\gamma}c^{\gamma}_{\alpha\beta}S(U_{\alpha}(x),A_{\beta}(x))\qquad\text{for $|\gamma|=k$}$
Then we find for $|\beta|=k$ that
(A.4) $S(U(x),\Delta_{h,y}A_{\beta}(x))=\Delta_{h,y}B_{\beta}(x)-S(\Delta_{h,y}U(x),A_{\beta}(x+hy))\\ -\Delta_{h,y}\sum_{\alpha\neq 0}c_{\alpha\beta}S(U_{\alpha}(x),A_{\beta-\alpha}(x))$
where the right hand side is continuous in $h$. As before, we find that that the left hand side has the limit $S(U,y\cdot\partial A_{k})(x)$ which gives that $A_{k}(x)\in C^{1}$ so $A(x)\in C^{k+1}$. By induction we obtain the result in the $C^{\infty}$ case.
If $U(x)$ and $B(x)$ are analytic, then they have holomorphic extensions in a complex neighborhood of $\mathbf{R}^{n}$. We then find that $S(U(z),\partial_{\bar{z}_{j}}A(z))=\partial_{\bar{z}_{j}}B(z)-S(\partial_{\bar{z}_{j}}U(z),A(z))=0$ in that neighborhood $\forall\,j$, which gives the analyticity of $A(x)$ by uniqueness. ∎
Appendix B Tame Fréchet spaces and the Nash–Moser theorem
In this section, we will recall the Nash–Moser Theorem and the theory of tame Fréchet spaces and tame maps from Sections II and III in [5]. Recall that a Fréchet space is a convex and complete vector space with topology defined by a countable number of seminorms, thus it is a complete metrizable vector space. But for the Nash–Moser Theorem one needs Fréchet spaces with a more refined structure.
Definition B.1.
A grading on a Fréchet space is a collection of seminorms $\{\|\cdot\|_{n},\ n\in\mathbf{N}\}$ that defines the topology and is increasing: $\|f\|_{j}\leq\|f\|_{k}$ when $j\leq k$.
Of course, every Fréchet space has a grading, for example by adding to each seminorm all the lower ones. Examples are $C^{\infty}$ functions and vector bundles with $C^{k}$ norms. Closed subspaces of graded spaces are graded with the induced norms, and Cartesian products $F\times G$ can be graded with the norms $\|(f,g)\|_{n}=\|f\|_{n}+\|g\|_{n}$.
Definition B.2.
Two gradings $\|\cdot\|_{n}$ and $\|\cdot\|^{\prime}_{n}$ are tamely equivalent of degree $r$ and base $b$ if
(B.1) $\|f\|_{n}\leq C_{n}\|f\|^{\prime}_{n+r}\quad\text{and}\quad\|f\|^{\prime}_{n}\leq C^{\prime}_{n}\|f\|_{n+r}$
for any $n\geq b$. If this holds for any $n$ then they are equivalent of degree $r$.
For example, if $X$ is a compact manifold, then the $C^{k}$ norms, Hölder norms and Sobolev norms are tamely equivalent.
Definition B.3.
A linear map $L:\ F\mapsto G$ of a graded space into another is tame, if it satisfies a tame estimate of degree $r$ and base $b$:
(B.2) $\|Lf\|_{n}\leq C_{n}\|f\|_{n+r}$
for any $n\geq b$. If (B.2) holds for any $n$ then $L$ is tame of degree $r$.
Example B.4.
A $C^{\infty}$ linear partial differential operator $L$ of order $m$ on $C^{\infty}$ is a tame of degree $m$ since $\|Lf\|_{n}\leq C_{n}\|f\|_{n+m}$ for the $C^{k}$ norms for any $n$. In particular, multiplication with $C^{\infty}$ functions are tame of degree 0. Solutions to invertible linear systems of smooth equations give tame maps of degree 0 by Leibniz’ rule. Similarly, for any $m\in\mathbf{N}$ and $C^{\infty}$ system ${\mathbb{M}}(t,x)$, we find that the mapping
$(Q(t,x),\mathbb{R}(t,x))\mapsto G(t,x)=Q(t,x)(t^{m}\operatorname{Id}_{n}+{\mathbb{M}}(t,x))+\mathbb{R}(t,x)$
is tame of degree 0.
Example B.5.
The linear mapping $L:C^{\infty}[0,1]\mapsto C^{\infty}[-1,1]$ given by $Lf(x)=f(x^{2})$ is tame with $\|Lf\|_{n}\leq C_{n}\|f\|_{n}$. The image of $L$ is the closed subspace of even functions $C_{S}^{\infty}[-1,1]$. But the inverse $L^{-1}:\ C_{S}^{\infty}[-1,1]\mapsto C^{\infty}[0,1]$ is not tame, since the $2n$:th Taylor coefficient of $Lf=g$ is the $n$:th Taylor coefficient of $f=L^{-1}g$ so the best estimate is $\|L^{-1}f\|_{n}\leq C_{n}\|f\|_{2n}$.
Compositions of linear tame maps are linear and tame, and $L$ is a tame isomorphism if $L$ is linear isomorphism such that $L$ and $L^{-1}$ are tame.
Definition B.6.
Let $F$ and $G$ be graded spaces. Then we say that $F$ is a tame direct summand of $G$ if we can find tame linear maps $L:\ F\mapsto G$ and $M:\ G\mapsto F$ so that the composition ML:FFML:\ $\mapsto F$ is the identity: $F\overset{L}{\mapsto}G\overset{M}{\mapsto}F$.
Definition B.7.
We say that a graded space is tame if it is a tame direct summand of a space $\Sigma(B)$ of exponentially decreasing sequences in some Banach space $B$.
A Cartesian product of two tame spaces is tame.
Theorem B.8.
The space $C^{\infty}(X)$ is tame if $X$ is a compact manifold (with or without boundary), and if $V$ is a vector bundle over $X$ then the sections of the vector bundle $C^{\infty}(X,V)$ is also tame, since it can be written as a tame direct summand of a tame space.
Theorem B.9.
If $X$ is a compact manifold with boundary, then $C^{\infty}_{0}(X)$ and $C^{\infty}_{0}(X,V)$ are tame.
Next, we shall look at nonlinear maps.
Definition B.10.
Let $F$ and $G$ be graded spaces and $P:\ F\supseteq U\mapsto G$ a nonlinear map, then we say that $P$ satisfies a tame estimate of degree $r$ and base $b$ if
(B.3) $\|P(f)\|_{n}\leq C_{n}(1+\|f\|_{n+r})$
for all $f\in U$ and all $n\geq b$. If (B.3) holds for any $n$ then $P$ is tame of degree $r$. We say that $P$ is a tame map if $P$ is defined on an open set, is continuous and satisfies a tame estimate in a neighborhood of any point.
As before, if (B.3) holds for any $n$ then we say $P$ satisfies a tame estimate of degree $r$.
Theorem B.11.
A composition of tame maps is tame.
The following is the Nash–Moser theorem that we shall use, which is Theorem 1.7.1 in section III of [5].
Theorem B.12.
Let $F$ and $G$ be tame spaces and $P:\ F\supseteq U\mapsto G$ a $C^{\infty}$ tame map of the open set $U$. If the equation for the derivative $DP(f)h=k$ has a unique solution $h=VP(f)k$ for any $f\in U$ and any $k$, and the family of inverses $VP(f):\ U\times G\mapsto F$ are tame maps, then $P$ is locally invertible and each local inverse is a $C^{\infty}$ tame map.
If the derivative $DP(f)h=k$ has a solution $h=VP(f)k$ for any $f\in U$ and any $k$, and the family of right inverses $VP(f):\ U\times G\mapsto F$ are tame maps, then $P$ is locally surjective and each local right inverse is a $C^{\infty}$ tame map.
If the derivative $DP(f)$ is injective for any $f\in U$ and the family of left inverses $VP(f):\ U\times G\mapsto F$ are tame maps, then $P$ is locally injective and each local left inverse is a $C^{\infty}$ tame map.
Remark B.13.
The local inverse $P^{-1}$ of $P$ in Theorem B.12 inherits the properties of the linear inverse VPVP, since $VP$ is used in the construction of $P^{-1}$, see Theorem 1.7.1 in section III of [5]. For example, if VPVP depends $C^{\infty}$ on some parameters then $P^{-1}$ also depends $C^{\infty}$ on these.
Example B.14.
It follows from Proposition 4.2 that if ${\mathbb{M}}(t)$ is a smooth symmetric $N\times N$ system valued function such that ${\mathbb{M}}(0)=0$ and $\partial_{t}^{m}{\mathbb{M}}\equiv 0$ then the mapping of smooth $N\times N$ system valued functions
(B.4) $C_{0}^{\infty}(\mathbf{R})\times C_{0}^{\infty}(\mathbf{R})\ni(Q(t),{\mathbb{M}}(t))\mapsto\\ G(t)=P(t,{\mathbb{M}}(t))Q(t)+R(t)\in C_{0}^{\infty}(\mathbf{R})\times C_{0}^{\infty}(\mathbf{R})$
is a $C^{\infty}$ tame surjective map of degree 0 with a local $C^{\infty}$ tame right inverse of degree $3+m$.
The proofs of these results and more examples and counterexamples can be found in sections II and III in [5].
Acknowledgment.
I would like to thank Gregory Berkolaiko at Texas A&M, who asked me about symmetric preparation theorems like (1.5) which started this project. We then worked on the formal power series solution in the first order case, see [4, Section 2].